OpenStudy (anonymous):

Which of the following pairs of slopes are of perpendicular lines? 1/2 and 2 3/4 and -4/3 -2/5 and -5/2

OpenStudy (anonymous):

how do you know if slopes are perpendicular?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

if one slope is \[\frac{ a }{ b }\] then a perpendicular slope is \[\frac{ -b }{ a }\]

OpenStudy (anonymous):

two slopes are perpendicular if m1 = -1/m2

OpenStudy (anonymous):

so it could be b and c?

OpenStudy (anonymous):

look at the negative signs in C..... aint gonna work

OpenStudy (anonymous):

ok thank you